OpenStudy (anonymous):

write a variable expression for a length of time in minutes if you know the number s of seconds.

OpenStudy (shubhamsrg):

t = s/60 ?? where s= seconds and t = time in minutes

OpenStudy (unklerhaukus):

\[t[\text {s}]\qquad\qquad\qquad60[\sec]=1[\text {min}]\]\[=t[\text {s}]\times1\qquad\qquad\qquad1=\frac{[\text {min}]}{60[\sec]}\]\[=t[\text {s}]\times\frac{[\text{min}]}{60[\text{s}]}\]\[t[\text s]=\frac{t[\text{min}]}{60}\] \[60t[\text s]={t[\text{min}]}\]